ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60,62 Lica , Dionis, University of Agronomical Sciences , Romania χ^{2} Test for Markov Chains. The Communication is a study of statistics having transformation probabilities depending on an unknown parameter. We demonstrate a teorem which generalizes the χ^{2} test for the Markov chains. We make statements which show that, if some conditions regarding the matrix of transformation are fultilled then the repartition of the researched statistics respects the normal rule. References: R. G. Chase, On the chi squarte test when the parameters are estimaded independety of the sample, Amer. Stat. vol. 67, Nr. 339 (1972), 609 - 611. C. M. Pomohaci, Verifying statistical hypotheses by χ^{2} test for Markov chains, Biotehnologii, ser. F, Nr.1 (1996), 109 - 115.